OpenStudy (prince1342):

Solve −2x^2 +3x − 9 = 0

OpenStudy (prince1342):

x equals quantity of 3 plus or minus 3i square root of 7 all over 4 x equals quantity of 3 plus or minus 9i square root of 7 all over 4 x equals quantity of negative 3 plus or minus 3i square root of 7 all over 4 x equals quantity of negative 3 plus or minus 9i square root of 7 all over 4

OpenStudy (koikkara):

well, can u please say how it looks like.... to you ?

OpenStudy (prince1342):

what do you mean?

OpenStudy (koikkara):

you can use your textbook and say what quadratic form it is in....

OpenStudy (prince1342):

its in standard form

OpenStudy (koikkara):

click this > http://www.tiger-algebra.com/drill/2x~2_3x-9=0/

OpenStudy (prince1342):

x=−b±b2−4ac√2awhenax2+bx+c=0

OpenStudy (koikkara):

go through the link, it will help u understand.

OpenStudy (koikkara):

did u get... anything like... 2x^2 + 3x - 9 = 0 2x^2 - 3x + 9 = 0 x = [3 ± √((-3)^2 - 4 * 2 * 9)] / (2 * 2) x = (3 ± √-63) / 4 x = (3 ± 3i√7) / 4

OpenStudy (prince1342):

no i got nothing like my answer choices